6. A helium balloon is designed to lift a load of 400 kg. Find the radius of this balloon?
media%2F2ab%2F2ab741a0-b8d8-4c38-8512-85
Answer #1

Since density of Helium is less than the density of air , force of buoyancy will be more than the gravitational force on the ballon.

\rho_{air}Vg-\rho_{He}Vg=400g

\rho_{air}V-\rho_{He}V=400

(\rho_{air}-\rho_{He})\frac{4}{3}\pi R^3=400

R=\left ( \frac{400}{(\rho_{air}-\rho_{He})\frac{4}{3}\pi} \right )^{1/3}

R=\left ( \frac{400}{(1.29-0.179)\frac{4}{3}\pi} \right )^{1/3}=4.41\,m

Radius of balloon is \boxed{R=4.41\,m}